# Zeyu Shen1    Basileal Imana1    Tong Wu1    Chong Xiang2    Prateek Mittal1    Aleksandra Korolova1

### v1

@@ Abstract
Abstract. Retrieval-Augmented Generation (RAG) enhances Large Language Models by grounding their outputs in external documents. These systems, however, remain vulnerable to attacks on the retrieval corpus, such as prompt injection. RAG-based search systems (e.g., Google’s Search AI Overview) present an interesting setting for studying and protecting against such threats, as defense algorithms can benefit from built-in reliability signals—like document ranking—and represent a non-LLM challenge for the adversary due to decades of work to thwart SEO.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a powerful solution to overcome the limitations of Large Language Models (LLMs) that rely solely on fixed, parametric knowledge that may be incomplete or outdated [18, 34, 52, 6]. By retrieving relevant documents from an external corpus and incorporating them into a model’s input, RAG enables more up-to-date, and contextually grounded responses. One prominent application of the RAG paradigm is its use in search engines augmented with language models. In these systems, a web search engine acts as the retriever, identifying documents relevant to the user’s query. The retrieved content is then passed to a language model (“the LLM”), which generates a final response grounded in the retrieved documents. Notable examples include Bing Chat [5], Perplexity AI [54], ChatGPT’s Search [51], and Google Search with AI Overviews [16].
RAG-based Web search is vulnerable.
@@
Despite the promise of RAG systems, they are vulnerable to adversarial attacks that undermine the quality of the generated responses. The retrieval corpus, in particular, is vulnerable to adversarial attacks such as corpus poisoning [70] and prompt injection attacks [17] that can manipulate the LLM to generate incorrect or even malicious responses [45]. Additional robustness challenges that undermine the effectiveness of RAG systems (but are not the focus of this work) include presence of noisy [67, 10], contradictory [64, 7], or unreliable documents [23, 29, 45].
Existing defenses have limited practicality.
@@
Existing frameworks proposed to enhance robustness in RAG exhibit limitations that hinder their practical application. In particular, RobustRAG [61], a major existing RAG framework aimed at providing adversarial robustness, suffers from limited performance in benign (no-attack) scenarios and struggles in complex generation tasks. It employs the natural strategy of majority voting over retrieved documents to mitigate the impact of adversarially manipulated contents. However, RobustRAG’s implementation of this strategy — based on either keywords or next-token probabilities — necessarily comes with significant information loss. In fact, defining a meaningful majority vote over free-form natural language is far from trivial [42].
Opportunity: leveraging document reliability metrics.
@@
RAG-based search presents a particularly interesting adversarial setting because it includes built-in reliability signals—such as document ranking—that are difficult for adversaries to circumvent. For example, for a query “best selling sedan in the US,” a malicious actor who aims to have their own product recommended in the LLM’s response must first successfully appear among the top search results [45]. This requires overcoming highly sophisticated defenses against search engine optimization (SEO) attacks that have been refined over more than two decades to prioritize credible and high-authority sources [26, 1, 56].
@@
Current RAG defenses overlook such reliability signals and treat retrieved documents as an unordered set [61, 60, 58, 69]. This oversight is a missed opportunity to layer complementary safeguards as part of a defense-in-depth approach [57, 21]. Signals such as search engine ranking generally correlate with information quality and trustworthiness. Lower-ranked documents, for instance, may be inherently noisier and also represent easier targets for retrieval corruption by adversaries.
Our contributions.
@@
We introduce ReliabilityRAG, a novel framework designed to make RAG-based systems robust. We present a surprisingly effective strategy of finding a “consistent majority” over a set of retrieved documents by taking a graph-theoretic perspective. Moreover, our approach explicitly incorporates reliability signals from the retriever—whether in the form of document rank or explicit reliability scores—to guide more robust generation. Our approach demonstrates superior adversarial robustness, effectively maintains utility on benign inputs, and excels in complex tasks requiring more extensive outputs (e.g., long-form generation), representing a significant step towards more effective and provably robust defenses against retrieval corruption in RAG.
@@
Our first contribution is a document-selection algorithm that identifies a “consistent majority” among retrieved documents by finding the Maximum Independent Set (MIS) [35] on a “contradiction graph.” In this graph, vertices represent the retrieved documents, and edges connect pairs determined to be contradictory by a Natural Language Inference (NLI) model [38]. [^2] A Natural Language Inference (NLI) model determines the logical relationship between two text statements (a ”premise” and a ”hypothesis”) by classifying their relations as neutral, entailment, or contradiction. In other words, we aim to identify the largest possible subset of mutually consistent documents. Crucially, when multiple such sets exist, our method prioritizes the set containing higher-ranked (i.e. more reliable) documents. This approach provides provable robustness guarantees against a bounded number of adversarial corruptions under natural assumptions about the attack and NLI model performance.
@@
Recognizing that finding the exact MIS is computationally expensive (as MIS takes exponential time) and thus may be infeasible for a large number of retrieved documents, especially in applications such as search where individuals expect answers quickly, our second contribution is to propose a general weighted sample and aggregate framework [48, 44] for this setting. This framework efficiently handles large retrieval sets by sampling smaller subsets based on document weights (reflecting reliability) and aggregating the results. It can be combined with various aggregation mechanisms, including our MIS approach, preserving some robustness guarantees while scaling effectively.
@@
To summarize, our key contributions are: (i) formalize the problem of RAG incorporating document reliability signals in Section 2; (ii) introduce a MIS-based algorithm for robust, reliability-aware document selection in Section 3; (iii) propose a general, scalable weighted sample and aggregate framework that preserves robustness for large document sets in Section 4; (iv) provide provable robustness guarantees for both approaches under natural assumptions in Section 3.2 and Appendix B.3; (v) empirically validate our methods, showing ReliabilityRAG achieves superior robustness to adversarial attacks and maintains high benign accuracy in Section 5. A detailed discussion of related works, including detailed comparisons with prior approaches for RAG robustness, is presented in Appendix A. In Appendix 5.6, we provide an empirical, end-to-end latency breakdown of our methods and practical speed-up tips for runtime-sensitive deployment settings.
@@ 2.1 RAG with Ordinal and Cardinal Reliability
Given a query $q$, the retriever returns an ordered list of $k$ documents $D=(x_{1},x_{2},\dots,x_{k})$. We will consistently use the convention that $x_{1}$ is the highest-ranked (most reliable) document and $x_{k}$ is the lowest-ranked (least reliable) document. Therefore, the higher the index, the lower the reliability.
A key distinction is whether the retriever supplies ordinal or cardinal reliability information about the retrieved documents; our defense can take advantage of either form. In the ordinal-reliability (rank-only) setting, we observe only the ordering $x_{1}\succeq x_{2}\succeq\dots\succeq x_{k}$, interpreted as $x_{1}$ is at least as reliable as $x_{2}$,” and so on. In the cardinal-reliability (rank + weight) setting, each document additionally carries a non-negative weight $w(x_{i})\in[0,1]$ with $w(x_{1})\geq w(x_{2})\geq\dots\geq w(x_{k})$, capturing graded reliability, with higher weight corresponding to more reliability (e.g. PageRank [53], citation count [12], or a learned reliability score [23, 59]).
@@ 2.2 Threat Model: Corrupted Documents in Retrieval
We consider a targeted attack on RAG-based search systems like Google’s AI Overview or ChatGPT Search. We focus on attacks that can corrupt some of the documents retrieved by the RAG-based search system. Attacks that directly target the operational infrastructure of the system provider (e.g., exploiting software vulnerabilities in Google or Microsoft machines) are out of scope. Motivated by extensive work in information retrieval that enables effective prioritization of authoritative and credible sources [26, 1, 56, 45] and withstands SEO attacks, our threat model captures the relative difficulty for the adversary to poison higher-ranked documents compared to lower-ranked ones.
Adversary Goal.
@@
We focus on attacks where the adversary’s objective is to induce a specific output in the LLM, such as inclusion of their own product in the AI overview. We assume that the attacker is able to inject $k^{\prime}$ documents into the $k$ documents returned by the retriever in response to a query, but injecting documents into the higher-ranked or higher-weighted positions is more difficult than doing so for the lower-ranked or lower-weighted ones. Formally, the attacker selects a subset $S\subseteq\{1,\ldots,k\}$ of size $k^{\prime}$, replacing those documents with arbitrary content, while the remaining documents are left unchanged. We assume bounded corruption, i.e. $k^{\prime}\ll k$, otherwise a robust and accurate defense is fundamentally impossible.
@@
For our empirical evaluations in Section 5, we specifically consider two corruption strategies commonly studied in prior work: (i) Corpus poisoning [70], which inserts false or misleading factual statements, and (ii) Prompt injection [17], which embeds jailbreak or control prompts to steer generation of the LLM.
Defense Goal.
@@
The objective of our defense is to sub-select from $D$ those documents that were not corrupted by the adversary, and pass only those to the LLM. If successful, our defense would thwart the attacker’s ability to produce a specific, malicious output from the LLM. Thus, in our first theoretical result in Section 3, we will be focused on computing the success probability of our sub-selection algorithm, i.e., the probability $\lambda$ that the subset of documents we pass to the LLM contains only benign documents. We call such framework $\lambda$-robust.
@@
Our approach forms a natural defense-in-depth: an attacker would first need to overcome sophisticated defenses built into retrievers (e.g., to thwart SEO attacks) to be included in the set of retrieved documents, and then bypass our reliability-aware filtering mechanism to be included in the subset of documents we pass to the LLM.
@@
Our defense objective yields a natural trade-off between robustness and utility. A perfect defense would filter out all malicious documents while maximally maintaining all benign documents. However, if the defense incurs false positives and additionally filters out some benign documents, it may have an impact on system utility. On the other hand, even if the defense incurs false negatives and leaves some malicious documents, the ultimate answer may still not be the one the adversary targeted. Therefore, we empirically evaluate the accuracy (utility) of the defense with aid of LLM-as-a-judge [68] in both benign scenarios and under attack in Section 5.
@@ 3 Ordinal-Reliability Setting: MIS-Based Algorithm
An independent set of an undirected graph $G=(V,E)$ is a subset $S\subseteq V$ such that no two vertices in $S$ are adjacent. MIS is a largest such subset. Although finding any MIS in the graph is NP-hard and, under the Exponential Time Hypothesis, requires $2^{\Omega(|V|)}$ time in the worst case [35], for graphs with up to a few dozen vertices the exact MIS can be found with brute-force in milliseconds. As RAG pipelines currently retrieve only $k\leq 20$ documents [43, 58, 60], finding MIS on a contradiction graph over retrieved documents is computationally practical. When $k$ grows larger, we resort to a sampling-based approach (Section 4), which preserves robustness guarantees with high probability while scaling to larger retrieval sets.
@@ 3.1 Rank-Aware Selection via MIS
In this section, we detail the procedure of rank-aware document selection via MIS. We first construct a contradiction graph with three steps: (i) Isolated Answering: For each of the top-$k$ retrieved documents $x_{i}$ (ranked $x_{1}\succeq x_{2}\succeq\cdots\succeq x_{k}$), the LLM is queried with the original query $q$ and the individual documents $q]+[x_{i}]$ to generate an isolated answer $y_{i}$. (ii) Contradiction Testing: An NLI model tests every pair of answers $(y_{i},y_{j})$; if the probability of the contradiction label exceeds a threshold $\beta$ (following  [55], we set $\beta=0.5$ in all experiments), the pair is deemed contradictory. (iii) Graph Encoding: The pairwise contradiction results are encoded into an undirected graph $G=(V,E)$, where the vertices $V$ represent the relevant documents (inheriting their retrieval rank), and an edge $(i,j)$ exists in $E$ iff the corresponding answers $(y_{i},y_{j})$ were deemed contradictory.
@@ 16        $S^{\star}\leftarrow S$
After $G$ is constructed, we enumerate all $2^{|V|}$ subsets (brute-force with bit-masking suffices for $|V|\leq k\leq 20$) and keep those that are independent. Among these we choose:
$S^{\star}\;=\;\arg\max\nolimits_{S\;\text{independent}}\bigl(\,|S|,\,-\mathrm{lex}(S)\bigr),$
where $\mathrm{lex}(S)$ is the lexicographic ordering of the vertex indices. Thus, intuitively, our rank-aware selection aims to maximize robustness through the search of maximal non-contradictory sets and to incorporate ordinal reliability signals through the choice among those sets. Ultimately, we query the LLM with $q]+\{x_{i}\}_{i\in S^{*}}$. We present the pseudocode for the full pipeline in Algorithm 1; $LLM(q,S)$ denotes the answer generated by the LLM prompted with query $q$ and set of documents $S$; $NLI(y_{i},y_{j})$ denotes the probability of the contradiction label judged by the NLI model for answers $y_{i}$ and $y_{j}$.

### v2

# ReliabilityRAG: Effective and Provably Robust Defense for RAG-based Web-Search

@@ Abstract
Retrieval-Augmented Generation (RAG) enhances Large Language Models by grounding their outputs in external documents. These systems, however, remain vulnerable to attacks on the retrieval corpus, such as prompt injection. RAG-based search systems (e.g., Google’s Search AI Overview) present an interesting setting for studying and protecting against such threats, as defense algorithms can benefit from built-in reliability signals—like document ranking—and represent a non-LLM challenge for the adversary due to decades of work to thwart SEO.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a powerful solution to overcome the limitations of Large Language Models (LLMs) that rely solely on fixed, parametric knowledge that may be incomplete or outdated Guu et al. (2020); Lewis et al. (2020); OpenAI et al. (2024); Brown et al. (2020). By retrieving relevant documents from an external corpus and incorporating them into a model’s input, RAG enables more up-to-date, and contextually grounded responses. One prominent application of the RAG paradigm is its use in search engines augmented with language models. In these systems, a web search engine acts as the retriever, identifying documents relevant to the user’s query. The retrieved content is then passed to a language model (“the LLM”), which generates a final response grounded in the retrieved documents. Notable examples include Bing Chat Bing (2025), Perplexity AI Perplexity (2025), ChatGPT’s Search OpenAI (2024c), and Google Search with AI Overviews Google (2025).
@@
RAG-based Web search is vulnerable. Despite the promise of RAG systems, they are vulnerable to adversarial attacks that undermine the quality of the generated responses. The retrieval corpus, in particular, is vulnerable to adversarial attacks such as corpus poisoning Zou et al. (2025) and prompt injection attacks Greshake et al. (2023) that can manipulate the LLM to generate incorrect or even malicious responses Nestaas et al. (2024). Additional robustness challenges that undermine the effectiveness of RAG systems (but are not the focus of this work) include presence of noisy Zhao et al. (2024); Cuconasu et al. (2024), contradictory Xu et al. (2024); Chen et al. (2022), or unreliable documents Hwang et al. (2025); Kumar and Lakkaraju (2024); Nestaas et al. (2024).
@@
Existing defenses have limited practicality. Existing frameworks proposed to enhance robustness in RAG exhibit limitations that hinder their practical application. In particular, RobustRAG Xiang et al. (2024a), a major existing RAG framework aimed at providing adversarial robustness, suffers from limited performance in benign (no-attack) scenarios and struggles in complex generation tasks. It employs the natural strategy of majority voting over retrieved documents to mitigate the impact of adversarially manipulated contents. However, RobustRAG’s implementation of this strategy — based on either keywords or next-token probabilities — necessarily comes with significant information loss. In fact, defining a meaningful majority vote over free-form natural language is far from trivial Mostafazadeh Davani et al. (2022).
@@
Opportunity: leveraging document reliability metrics. RAG-based search presents a particularly interesting adversarial setting because it includes built-in reliability signals—such as document ranking—that are difficult for adversaries to circumvent. For example, for a query “best selling sedan in the US,” a malicious actor who aims to have their own product recommended in the LLM’s response must first successfully appear among the top search results Nestaas et al. (2024). This requires overcoming highly sophisticated defenses against search engine optimization (SEO) attacks that have been refined over more than two decades to prioritize credible and high-authority sources Joachims et al. (2005); Ankalkoti (2017); Shahzad et al. (2020).
@@
Current RAG defenses overlook such reliability signals and treat retrieved documents as an unordered set Xiang et al. (2024a); Wei et al. (2025); Wang et al. (2024a); Zhou et al. (2025). This oversight is a missed opportunity to layer complementary safeguards as part of a defense-in-depth approach Stytz (2004); Holmberg (2017). Signals such as search engine ranking generally correlate with information quality and trustworthiness. Lower-ranked documents, for instance, may be inherently noisier and also represent easier targets for retrieval corruption by adversaries.
@@
Our contributions. We introduce ReliabilityRAG, a novel framework designed to make RAG-based systems robust. We present a surprisingly effective strategy of finding a “consistent majority” over a set of retrieved documents by taking a graph-theoretic perspective. Moreover, our approach explicitly incorporates reliability signals from the retriever—whether in the form of document rank or explicit reliability scores—to guide more robust generation. Our approach demonstrates superior adversarial robustness, effectively maintains utility on benign inputs, and excels in complex tasks requiring more extensive outputs (e.g., long-form generation), representing a significant step towards more effective and provably robust defenses against retrieval corruption in RAG.
@@
Our first contribution is a document-selection algorithm that identifies a “consistent majority” among retrieved documents by finding the Maximum Independent Set (MIS) Liu (1992) on a “contradiction graph.” In this graph, vertices represent the retrieved documents, and edges connect pairs determined to be contradictory by a Natural Language Inference (NLI) model MacCartney (2009). [^1] A Natural Language Inference (NLI) model determines the logical relationship between two text statements (a ”premise” and a ”hypothesis”) by classifying their relations as neutral, entailment, or contradiction. In other words, we aim to identify the largest possible subset of mutually consistent documents. Crucially, when multiple such sets exist, our method prioritizes the set containing higher-ranked (i.e. more reliable) documents. This approach provides provable robustness guarantees against a bounded number of adversarial corruptions under natural assumptions about the attack and NLI model performance.
@@
Recognizing that finding the exact MIS is computationally expensive (as MIS takes exponential time) and thus may be infeasible for a large number of retrieved documents, especially in applications such as search where individuals expect answers quickly, our second contribution is to propose a general weighted sample and aggregate framework Nissim et al. (2007); Near and Abuah (2024) for this setting. This framework efficiently handles large retrieval sets by sampling smaller subsets based on document weights (reflecting reliability) and aggregating the results. It can be combined with various aggregation mechanisms, including our MIS approach, preserving some robustness guarantees while scaling effectively.
@@
To summarize, our key contributions are: (i) formalize the problem of RAG incorporating document reliability signals in Section 2; (ii) introduce a MIS-based algorithm for robust, reliability-aware document selection in Section 3; (iii) propose a general, scalable weighted sample and aggregate framework that preserves robustness for large document sets in Section 4; (iv) provide provable robustness guarantees for both approaches under natural assumptions in Section 3.2 and Appendix B.3; (v) empirically validate our methods, showing ReliabilityRAG achieves superior robustness to adversarial attacks and maintains high benign accuracy in Section 5. A detailed discussion of related works, including detailed comparisons with prior approaches for RAG robustness, is presented in Appendix A. In Appendix D.2, we provide an empirical, end-to-end latency breakdown of our methods and practical speed-up tips for runtime-sensitive deployment settings.
@@ 2.1 RAG with Ordinal and Cardinal Reliability
Given a query $q$, the retriever returns an ordered list of $k$ documents $D=(x_{1},x_{2},\dots,x_{k})$. We will consistently use the convention that $x_{1}$ is the highest-ranked (most reliable) document and $x_{k}$ is the lowest-ranked (least reliable) document. When we refer to “higher-ranked” documents, we mean those closer to the front of the list (smaller index, greater reliability), and “lower-ranked” documents are those closer to the end of the list (larger index, lower reliability).
A key distinction is whether the retriever supplies ordinal or cardinal reliability information about the retrieved documents; our defense can take advantage of either form. In the ordinal-reliability (rank-only) setting, we observe only the ordering $x_{1}\succeq x_{2}\succeq\dots\succeq x_{k}$, interpreted as $x_{1}$ is at least as reliable as $x_{2}$,” and so on. In the cardinal-reliability (rank + weight) setting, each document additionally carries a non-negative weight $w(x_{i})\in[0,1]$ with $w(x_{1})\geq w(x_{2})\geq\dots\geq w(x_{k})$, capturing graded reliability, with higher weight corresponding to more reliability (e.g. PageRank Page et al. (1999), citation count Ding et al. (2009), or a learned reliability score Hwang et al. (2025); Wang et al. (2024b)).
@@ 2.2 Threat Model: Corrupted Documents in Retrieval
We consider a targeted attack on RAG-based search systems like Google’s AI Overview or ChatGPT Search. We focus on attacks that can corrupt some of the documents retrieved by the RAG-based search system. Attacks that directly target the operational infrastructure of the system provider (e.g., exploiting software vulnerabilities in Google or Microsoft machines) are out of scope. Motivated by extensive work in information retrieval that enables effective prioritization of authoritative and credible sources Joachims et al. (2005); Ankalkoti (2017); Shahzad et al. (2020); Nestaas et al. (2024) and withstands SEO attacks, our threat model captures the relative difficulty for the adversary to poison higher-ranked documents compared to lower-ranked ones.
@@
Adversary Goal. We focus on attacks where the adversary’s objective is to induce a specific output in the LLM, such as inclusion of their own product in the AI overview. We assume that the attacker is able to inject $k^{\prime}$ documents into the $k$ documents returned by the retriever in response to a query, but injecting documents into the higher-ranked or higher-weighted positions is more difficult than doing so for the lower-ranked or lower-weighted ones. Formally, the attacker selects a subset $S\subseteq\{1,\ldots,k\}$ of size $k^{\prime}$, replacing those documents with arbitrary content, while the remaining documents are left unchanged. We assume bounded corruption, i.e. $k^{\prime}\ll k$, otherwise a robust and accurate defense is fundamentally impossible.
@@
For our empirical evaluations in Section 5, we specifically consider two corruption strategies commonly studied in prior work: (i) Corpus poisoning Zou et al. (2025), which inserts false or misleading factual statements, and (ii) Prompt injection Greshake et al. (2023), which embeds jailbreak or control prompts to steer generation of the LLM.
@@
Defense Goal. The objective of our defense is to sub-select from $D$ those documents that were not corrupted by the adversary, and pass only those to the LLM. If successful, our defense would thwart the attacker’s ability to produce a specific, malicious output from the LLM. Thus, in our first theoretical result in Section 3, we will be focused on computing the success probability of our sub-selection algorithm, i.e., the probability $\lambda$ that the subset of documents we pass to the LLM contains only benign documents. We call such framework $\lambda$-robust.
@@
Our approach forms a natural defense-in-depth: an attacker would first need to overcome sophisticated defenses built into retrievers (E.g., to thwart SEO attacks), and then bypass our reliability-aware filtering mechanism.
@@
Our defense objective yields a natural trade-off between robustness and utility. A perfect defense would filter out all malicious documents while maximally maintaining all benign documents. However, if the defense incurs false positives and additionally filters out some benign documents, it may have an impact on system utility. On the other hand, even if the defense incurs false negatives and leaves some malicious documents, the ultimate answer may still not be the one the adversary targeted. Therefore, we empirically evaluate the accuracy (utility) of the defense with aid of LLM-as-a-judge Zheng et al. (2023) in both benign scenarios and under attack in Section 5.
@@ 3 Ordinal-Reliability Setting: MIS-Based Algorithm
An independent set of an undirected graph $G=(V,E)$ is a subset $S\subseteq V$ such that no two vertices in $S$ are adjacent. MIS is a largest such subset. Although finding any MIS in the graph is NP-hard and, under the Exponential Time Hypothesis, requires $2^{\Omega(|V|)}$ time in the worst case Liu (1992), for graphs with up to a few dozen vertices the exact MIS can be found with brute-force in milliseconds. As RAG pipelines currently retrieve only $k\leq 20$ documents MyScale (2024); Wang et al. (2024a); Wei et al. (2025), finding MIS on a contradiction graph over retrieved documents is computationally practical. When $k$ grows larger, we resort to a sampling-based approach (Section 4), which preserves robustness guarantees with high probability while scaling to larger retrieval sets.
@@ 3.1 Rank-Aware Selection via MIS
In this section, we detail the procedure of rank-aware document selection via MIS. We first construct a contradiction graph with three steps: (i) Isolated Answering: For each of the top-$k$ retrieved documents $x_{i}$ (ranked $x_{1}\succeq x_{2}\succeq\cdots\succeq x_{k}$), the LLM is queried with the original query $q$ and the individual documents $q]+[x_{i}]$ to generate an isolated answer $y_{i}$. (ii) Contradiction Testing: An NLI model tests every pair of answers $(y_{i},y_{j})$; if the probability of the contradiction label exceeds a threshold $\beta$ (following  Schuster et al. (2022), we set $\beta=0.5$ in all experiments), the pair is deemed contradictory. (iii) Graph Encoding: The pairwise contradiction results are encoded into an undirected graph $G=(V,E)$, where the vertices $V$ represent the relevant documents (inheriting their retrieval rank), and an edge $(i,j)$ exists in $E$ iff the corresponding answers $(y_{i},y_{j})$ were deemed contradictory.
@@ 16        $S^{\star}\leftarrow S$
After $G$ is constructed, we enumerate all $2^{|V|}$ subsets (brute-force with bit-masking suffices for $|V|\leq k\leq 20$) and keep those that are independent. Among these we choose $S^{\star}\;=\;\arg\max\nolimits_{S\;\text{independent}}\bigl(\,|S|,\,-\mathrm{lex}(S)\bigr),$ where $\mathrm{lex}(S)$ is the lexicographic ordering of the vertex indices. Thus, intuitively, our rank-aware selection aims to maximize robustness through the search of maximal non-contradictory sets and to incorporate ordinal reliability signals through the choice among those sets. Ultimately, we query the LLM with $q]+\{x_{i}\}_{i\in S^{*}}$. We present the pseudocode for the full pipeline in Algorithm 1; $LLM(q,S)$ denotes the answer generated by the LLM prompted with query $q$ and set of documents $S$; $NLI(y_{i},y_{j})$ denotes the probability of the contradiction label judged by the NLI model for answers $y_{i}$ and $y_{j}$.